OpenStudy (anonymous):

factor x^4+x^2+1(add and subtract x^2)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

i need some help with these

OpenStudy (phi):

I guess we should do what they suggest x^4+x^2+1 + x^2 - x^2 or x^4+2 x^2+1 - x^2 if we label x^2 as y , the first 3 terms give y^2 +2y + 1 which looks like a perfect square can you continue?

OpenStudy (anonymous):

if we subtract first we get-(x^2-x^4-x^2-1)

OpenStudy (phi):

No, we do not do that. start with \[ x^4+x^2+1 \] we can add zero to this expression and not change it \( x^2 - x^2 =0 \) \[ x^4+x^2+1 + x^2 - x^2\] or \[ x^4+2x^2+1 - x^2 \] now focus just on \[ x^4+2x^2+1 \] can you factor that ? (it is the same as y^2 + 2y + 1... which factors nicely)

OpenStudy (anonymous):

x(x^3+2x)(x+1)

OpenStudy (anonymous):

idk i'm messing up i'm bad at these

OpenStudy (phi):

ok. can you factor y^2 + 2y + 1 ? look at the last number, and list all pairs of factors for example, if the last number was 6 we would list: 1,6 2,3 but in this case we have 1, and all we can list is 1,1 now look at the pairs on the list (here we have only 1 choice) do they add up to the middle number (the 2 in 2y) if so, they are our factors. we write (y+1)(y+1)

OpenStudy (anonymous):

yes

OpenStudy (phi):

notice that if we rename x^2 to y we can do this \[ x^4 + 2 x^2 + 1\\(x^2)^2 + 2 x^2 + 1 \\ y^2 + 2y +1\\(y+1)(y+1)\\(y+1)^2\] now replace y with x^2 to get the answer \[ (x^2+1)^2\]

OpenStudy (phi):

now you have \[ x^4+2 x^2+1 - x^2 \\ (x^2+1)^2 - x^2\] that is a difference of squares use this rule \[ a^2 - b^2 = (a-b)(a+b) \] in this problem, a = (x^2+1) and b=x

OpenStudy (anonymous):

x^2+1^2-x^2=(x^2+1-x)(x^2+1+x)

OpenStudy (phi):

ok but you should put parens in: x^2+1^2-x^2 should be written (x^2+1)^2-x^2 but so far, so good. people write the polynomial in standard form.. x^2 -x +1 (the x's are listed according to their exponent) so the answer is (x^2-x+1)(x^2+x+1)